An archer shoots an arrow towards a 300g target that is sliding in her direction at a speed of 2.50m/s on a smooth frictionless surface. The 22.5g arrow is shot with speed of 35m/s and passes through the target without changing the mass of the target, which is stopped by the impact. Calculate the speed of the arrow after passing through the target?

1
Expert's answer
2022-05-11T13:53:05-0400

"u=\frac{m_1v_1-m_2v_2}{m_1}=30~\frac ms."